# Quantum variational calculus on a lattice

### v1

@@ 5.2. Conserved charge associated to global $U(1)$ symmetry
Another example of symmetries we have not considered so far with the base a lattice or a discrete Abelian group are vertical continuous symmetries as in Example 2.7 in the continuum case. Here we can copy everything from the continuum, as the symmetry only acts on the vertical 1-forms which form a Grassmann algebra. We consider a complex scalar field on a discrete Abelian group $X$ with inverse metric given by $\lambda^{a}$, modelled by the bundle $X\times\mathbb{C}\to X$, meaning that the jet bundle will now have coordinates $\bar{u},u$ and the respective derivatives. We choose the action
@@
with the Lagrangian $L=\left(\sum_{a}{1\over 4}\bar{u}_{a}u_{a}-{m^{2}\over 2}\bar{u}u\right)$. This system has a global $U(1)$-symmetry $\phi\mapsto e^{{\mathfrak{i}}\varphi}\phi$, $\phi^{*}\mapsto e^{-{\mathfrak{i}}\varphi}\phi^{*}$, which can be implemented by setting $\iota_{V}\mathrm{d}_{V}u_{I}={\mathfrak{i}}\varphi u_{I}$ and $\iota_{V}\mathrm{d}_{V}\bar{u}_{I}=-{\mathfrak{i}}\varphi\bar{u}_{I}$, with $\iota_{H}=0$. The EL and boundary form are
@@
$j_{U(1)}=-\iota_{V}\Theta=-\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}R_{a^{-1}}\left(\mu\lambda^{a}\right)\left(\bar{u}_{a^{-1}}u-u_{a^{-1}}\bar{u}\right)\mathrm{Vol}_{a}.$
@@
$j_{U(1)}[\phi,\phi^{*}]=-\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}R_{a^{-1}}\left(\mu\lambda^{a}\right)\left((\partial_{a^{-1}}\phi^{*})\phi-(\partial_{a^{-1}}\phi)\phi^{*}\right)\mathrm{Vol}_{a}.$
@@
$\mathrm{d}_{X}j_{U(1)}[\phi,\phi^{*}]=\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}\partial_{a}\left(R_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi^{*}\right)\phi-R_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi\right)\phi^{*}\right)\mathrm{Vol}$
$=\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}\left(-\partial_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi^{*}\right)R_{a}\phi+R_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi^{*}\right)\partial_{a}\phi+\partial_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi\right)R_{a}\phi^{*}-R_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi\right)\partial_{a}\phi^{*}\right)$
$=\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}\left(-2\mu m^{2}\phi^{*}R_{a}\phi+\left(\mu\lambda^{a}\partial_{a}\phi^{*}+2\mu m^{2}\phi^{*}\right)\partial_{a}\phi+2\mu m^{2}\phi R_{a}\phi^{*}-\left(\mu\lambda^{a}\partial_{a}\phi+2\mu m^{2}\phi\right)\partial_{a}\phi^{*}\right)$
=0=0= 0
@@ 5.3. Complex scalar field theory on $X$ with background gauge field
Computing this expression and assuming the same relation between the measure $\mu$, inverse metric $\lambda^{a}$ and braiding $\sigma$ as in Proposition 5.1, we recover exactly the operator in (21), so that the latter becomes the covariantized Klein-Gordon equation $(\Delta_{\alpha}+m^{2})\phi=0$. We have done this for $X$ an Abelian group but as with the end of Section 5.1 there is also a version on any graph.
@@ 6. Concluding remarks
We have presented what we believe to be the first derivation of field equations and exactly conserved Noether charges for classical mechanics and field theory on a lattice from an action. We viewed the lattice as an exact noncommutative geometry and used a theory of jet bundles from our previous work [20], albeit an easy case as the base coordinate algebra remains commutative and only differential forms (given by arrows of the lattice) non-commute with functions. The resulting conserved quantities are not obvious even for a lattice line (but can be easily checked in this case). Having a good foundation for classical field theory including conserved quantities is also an important step for the Hamiltonian form of quantum field theory as a complement to existing methods even on a lattice[23].
There are several further directions to be explored in further work. First of all, as suggested by Proposition 5.3, our methods should extend to any graph, not necessarily an Abelian group lattice (which is the case covered here). The explicit approach to jet bundles in [20] requires a suitable torsion free flat background connection $\nabla$, which can be solved even for some nonAbelian group Cayley graphs, but can also be avoided altogether when coming form more abstract methods as in the subsequent work [11]. It should also be straightforward to cover the hybrid case of ×$\mathbb{R}\times G$ where $\mathbb{R}$ is a continuous time and $G$ is a discrete group or graph.
@@
We also limited ourselves to trivial bundles. The jet formalism itself and Anderson’s approach can include nontrivial bundles $E$ as (finitely generated projective) bimodules over the coordinate algebra $A$ and our approach on the lattice should extend to this with more care. Finally, the current methodology (even for trivial bundles) works similarly for noncommutative $A$ provided there are nice properties for the exterior algebra, such as a global basis. The simplest case of $A$ noncommutative but with a central Grassmann algebra basis is straightforward and would apply to $A$ the fuzzy sphere, for example. Other examples where there is no problem to define the jet bundle[20] and where our methods should work would be to variational calculus and hence classical field theory on the Majid-Ruegg $\kappa$-Minkowski spacetime.
Finally, while we found the Anderson-Zuckerman approach[1, 27] particularly natural for noncommutative geometry, there are other variants and routes to physics in the case of a classical manifold, e.g. [12, 22] which may have aspects that extend to the noncommutative case. At the mathematical level, another approach to jets bundles is as Hopf algebroids, see [14] where examples are obtained by cotwist. Note that knowing the noncommutative analogue of sections of the jet bundle as in [20, 11, 14] is not enough to know what is the noncommutative analogue of $C(J^{\infty})$ and its differential calculus. Also, jets are dual to differential operators, see [16]. This is the natural coordinate-invariant notion of the Heisenberg algebra for a classical manifold and could give another route to the physics possibly bypassing jets entirely. Here, a Hopf algebroid of differential operators on an algebra $A$ with suitable calculus $(\Omega,\mathrm{d})$ was found in [13], and examples by twisting the classical algebra of differential operators in [26].

### v2

@@ 5.2. Conserved charge associated to global $U(1)$ symmetry
Another example of symmetries that we have not considered so far with the base a lattice or a discrete Abelian group are vertical continuous symmetries as in Example 2.7 in the continuum case. Here we can copy everything from the continuum, as the symmetry only acts on the vertical 1-forms which form a Grassmann algebra. We consider a complex scalar field on a discrete Abelian group $X$ with inverse metric given by $\lambda^{a}$, modelled by the bundle $X\times\mathbb{C}\to X$, meaning that the jet bundle will now have coordinates $\bar{u},u$ and the respective derivatives. We choose the action
@@
with the Lagrangian $L=\left(\sum_{a}{1\over 4}\lambda^{a}\bar{u}_{a}u_{a}-{m^{2}\over 2}\bar{u}u\right)$. This system has a global $U(1)$-symmetry $\phi\mapsto e^{{\mathfrak{i}}\varphi}\phi$, $\phi^{*}\mapsto e^{-{\mathfrak{i}}\varphi}\phi^{*}$, which can be implemented by setting $\iota_{V}\mathrm{d}_{V}u_{I}={\mathfrak{i}}\varphi u_{I}$ and $\iota_{V}\mathrm{d}_{V}\bar{u}_{I}=-{\mathfrak{i}}\varphi\bar{u}_{I}$, with $\iota_{H}=0$. The EL and boundary form are
@@
$j_{U(1)}=-\iota_{V}\Theta=\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}R_{a^{-1}}\left(\mu\lambda^{a}\right)\left(\bar{u}_{a^{-1}}u-u_{a^{-1}}\bar{u}\right)\mathrm{Vol}_{a}.$
@@
$j_{U(1)}[\phi,\phi^{*}]=\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}R_{a^{-1}}\left(\mu\lambda^{a}\right)\left((\partial_{a^{-1}}\phi^{*})\phi-(\partial_{a^{-1}}\phi)\phi^{*}\right)\mathrm{Vol}_{a}.$
@@
$\mathrm{d}_{X}j_{U(1)}[\phi,\phi^{*}]=-\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}\partial_{a}\left(R_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi^{*}\right)\phi-R_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi\right)\phi^{*}\right)\mathrm{Vol}$
$=-\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}\left(-\partial_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi^{*}\right)R_{a}\phi+R_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi^{*}\right)\partial_{a}\phi+\partial_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi\right)R_{a}\phi^{*}-R_{a^{-1}}\left(\mu\lambda^{a}\partial_{a}\phi\right)\partial_{a}\phi^{*}\right)$
$=-\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}\left(-2\mu m^{2}\phi^{*}R_{a}\phi+\left(\mu\lambda^{a}\partial_{a}\phi^{*}+2\mu m^{2}\phi^{*}\right)\partial_{a}\phi+2\mu m^{2}\phi R_{a}\phi^{*}-\left(\mu\lambda^{a}\partial_{a}\phi+2\mu m^{2}\phi\right)\partial_{a}\phi^{*}\right)$
$=0$
@@
If there is a preferred time direction then we write the lattice as $\mathbb{Z}\times\mathbb{Z}^{m-1}$. From Corollary 4.3, there is a conserved charge $Q_{U(1)}[\phi,\phi^{*}]$ associated to $j_{U(1)}$ and given by
$Q_{U(1)}[\phi,\phi^{*}]={{\mathfrak{i}}\varphi\over 4}\int_{\mathbb{Z}^{m-1}}$ $\left[\left(R_{t^{-1}}(\mu\lambda^{t})+\mu\lambda^{t^{-1}}\right)((\partial_{t^{-1}}\phi^{*})\phi-(\partial_{t^{-1}}\phi)\phi^{*})\right$.
$\left.+\mu\lambda^{t^{-1}}((\partial_{t^{-1}}\phi^{*})\partial_{t}\phi-(\partial_{t^{-1}}\phi)\partial_{t}\phi^{*})\right]\mathrm{Vol}_{\mathbb{Z}^{m-1}}.$
From this, we can read the $U(1)$-charge density $\rho$ and the current density as having the form
$J^{x}_{U(1)}[\phi,\phi^{*}]={{\mathfrak{i}}\varphi\over 4}$ $\left[\left(R_{x^{-1}}(\mu\lambda^{x})+\mu\lambda^{x^{-1}}\right)((\partial_{x^{-1}}\phi^{*})\phi-(\partial_{x^{-1}}\phi)\phi^{*})\right$.
$\left.+\mu\lambda^{x^{-1}}((\partial_{x^{-1}}\phi^{*})\partial_{x}\phi-(\partial_{x^{-1}}\phi)\partial_{x}\phi^{*})\right].$
In the continuum limit, we expect $\partial_{t^{\pm 1}}\phi\mapsto\pm\partial_{0}\phi$, $\partial_{x^{\pm 1}}\phi\mapsto\pm\partial_{i}\phi$ and similar for $\phi^{*}$, that $R_{t^{-1}}\mapsto\mathrm{id}$ and $\lambda^{t^{\pm 1}}\mapsto\mathfrak{g}^{00}$, $\lambda^{x^{\pm 1}}\mapsto\mathfrak{g}^{ii}$. Therefore in both $Q_{U(1)}[\phi,\phi^{*}]$ and $J^{x}_{U(1)}[\phi,\phi^{*}]$, the second line vanishes and the first recovers the expression from Example 2.7.
For $\lambda$ the Euclidean or Minkowski metric with $\lambda^{t}=\lambda^{t^{-1}}=1$, $\lambda^{x}=\lambda^{x^{-1}}=\pm 1$, we can redo the calculations for the conserved current and stress-energy tensor as in Proposition 4.1, now taking the doubling of the fields into account to find
$j=-\sum_{a,b\in\mathcal{C}}\epsilon^{b}\left(\frac{\lambda^{a}}{4}\left(\bar{u}_{a^{-1}}+\frac{1}{2}\bar{u}_{ba^{-1}}\right)u_{b}+\frac{\lambda^{a}}{4}\left(u_{a^{-1}}+\frac{1}{2}u_{ba^{-1}}\right)\bar{u}_{b}+\delta^{a}_{b}L\right)\mathrm{Vol}_{a},$
$T_{ab}=-\left(\frac{1}{4}\bar{u}_{a}u_{b}+\frac{1}{4}u_{a}\bar{u}_{b}+\frac{1}{8}\bar{u}_{ab}u_{b}+\frac{1}{8}u_{ab}\bar{u}_{b}+\frac{\delta^{a^{-1}}_{b}}{\lambda^{a}}L\right).$
In the case of a plane wave solution $\phi(t,x)=Ae^{-{\mathfrak{i}}(\omega t+\lambda^{x}\kappa x)}$ in (1+1)-dimensions these result in the energy and momentum density
$\mathcal{E}_{dens}[\phi,\phi^{*}]=|A|^{2}\sin^{2}(\omega),\quad\quad\mathcal{P}_{x,dens}[\phi,\phi^{*}]=\lambda^{x}|A|^{2}\sin(\kappa)\sin(\omega).$
which as in continuum field theory correspond to 2 times the space averages in equation (22) of the real scalar field, since we now have two fields $\phi,\phi^{*}$. Again, we find similar relations as in Equations (23), (24),  (27) and  (28), but now for the energy and momentum densities directly instead of for their averages.
The $U(1)$-charge and current densities for these solutions read
$\rho_{U(1)}[\phi,\phi^{*}]=|A|^{2}\varphi\sin(\omega)\cos(\omega),\quad\quad J^{x}_{U(1)}[\phi,\phi^{*}]=|A|^{2}\varphi\sin(\kappa)\cos(\kappa),$
which recover the classical values in Example 2.7 for $\omega,\kappa\ll 1$.
@@ 5.3. Complex scalar field theory on $X$ with background gauge field
Computing this expression and assuming the same relation between the measure $\mu$, inverse metric $\lambda^{a}$ and braiding $\sigma$ as in Proposition 5.1, we recover exactly the operator in (34), so that the latter becomes the covariantized Klein-Gordon equation $(\Delta_{\alpha}+m^{2})\phi=0$. We have done this for $X$ an Abelian group but as with the end of Section 5.1, there is also a version on any graph.
@@ 6. Concluding remarks
We have presented what we believe to be the first derivation of field equations and exactly conserved Noether charges for classical mechanics and field theory on a lattice from an action. We viewed the lattice as an exact noncommutative geometry and used a theory of jet bundles from our previous work [21], albeit an easy case as the base coordinate algebra remains commutative and only differential forms (given by arrows of the lattice) non-commute with functions. The resulting conserved quantities are not obvious even for a lattice line (but can be easily checked in this case). Having a good foundation for classical field theory including conserved quantities is also an important step for the Hamiltonian form of quantum field theory as a complement to existing functional integral methods even on a lattice [26].
There are several further directions to be explored in further work. First of all, as suggested by Proposition 5.3, our methods should extend to any graph, not necessarily an Abelian group lattice (which is the case covered here). The explicit approach to jet bundles in [21] requires a suitable torsion free flat background connection $\nabla$, which can be solved even for some non-Abelian group Cayley graphs, but can also be avoided altogether when coming form more abstract methods as in the subsequent work [11]. It should also be straightforward to cover the hybrid case of $\mathbb{R}\times G$ where $\mathbb{R}$ is a continuous time and $G$ is a discrete group or graph.
@@
We also limited ourselves to trivial bundles. The jet formalism itself and the Anderson-Zuckerman approach [1, 30] can include nontrivial bundles $E$ as (finitely generated projective) bimodules over the coordinate algebra $A$ and our approach on the lattice should extend to this with more care. Finally, the current methodology (even for trivial bundles) works similarly for noncommutative $A$ provided there are nice properties for the exterior algebra, such as a global basis. The simplest case of $A$ noncommutative but with a central Grassmann algebra basis is straightforward and would apply to $A$ the fuzzy sphere, for example. Other examples where there is no problem to define the jet bundle [21] and where our methods should work would be to variational calculus and hence classical field theory on the Majid-Ruegg $\kappa$-Minkowski spacetime.
Finally, while we found the Anderson-Zuckerman approach particularly natural for noncommutative geometry, there are other variants and routes to physics in the case of a classical manifold, e.g. [12, 24] which may have aspects that extend to the noncommutative case. At the mathematical level, another approach to jet bundles is as Hopf algebroids, see [15] where examples are obtained by cotwist. Note that knowing the noncommutative analogue of sections of the jet bundle as in [21, 11, 15] is not enough to know what is the noncommutative analogue of $C(J^{\infty})$ and its differential calculus. Also, jets are dual to differential operators, see [17], with the latter the natural coordinate-invariant notion of the Heisenberg algebra for a classical manifold. This could give another route to the physics possibly bypassing jets entirely. Here, a Hopf algebroid of differential operators on a possibly noncommutative algebra $A$ with suitable calculus $(\Omega,\mathrm{d})$ was found in [13], and examples by twisting the classical algebra of differential operators in [29].